OpenStudy (anonymous):

if |z+1|=z+2(1+i),find z?

OpenStudy (helder_edwin):

put \(z=a+ib\). do all the operations on each side and you will get a system of equations

OpenStudy (anonymous):

When you distribute the 2 to the (1+i) you get |z+1|=z+2+2i. Since 2i=-1, the equation becomes |z+1|=z+2-1 which would simplify to |z+1|=z+1

OpenStudy (primeralph):

Nope.

OpenStudy (primeralph):

2i is not -1.

OpenStudy (anonymous):

oh crap. youre right. sorry...

OpenStudy (primeralph):